# MAGICS I. The First Few Orbits Encode the Fate of Seed Massive Black Hole Pairs

### v1

@@ Abstract
keywords:
black hole physics – galaxies: kinematics and dynamics – galaxies: high-redshift – quasars: supermassive black holes – methods: numerical
pubyear: 2023
@@ 1 Introduction
Currently, our most accurate understanding of the MBH binary population and its relation with galaxy evolution comes from cosmological simulations (see e.g. Amaro-Seoane et al., 2023, for an overview). Cosmological simulations self-consistently follow the coevolution of MBHs with host galaxies, and contain rich information about the environments where MBH interactions and mergers take place (e.g. Salcido et al., 2016; Kelley et al., 2017; Tremmel et al., 2017; Volonteri et al., 2020; Katz et al., 2020; Chen et al., 2022). These environments include a wide range of scenarios from isolated dwarf galaxies to infalling satellites of a massive central galaxy in a cluster. However, realistic and large-volume modeling comes at the cost of limited resolutions, and at best they can follow MBH dynamics to $\sim{\rm kpc}$ scales. These simulations are also subject to the simplistic subgrid-seeding mechanism that only considers the higher end of the MBH seed mass (out of these, Dubois et al., 2014; Tremmel et al., 2015, probes relatively low MBH seeds at high-redshifts). To compensate for the resolution limit in large-volume simulations, cosmological “zoom-in” simulations are also applied to study the sinking behavior of specific merging systems (e.g. Pfister et al., 2019; Bortolas et al., 2020), but they are also subjected to high-computational cost and low flexibilities in configuration and subgrid models. Only a few galaxies and the MBHs can be studied at a time.
@@
To truly bridge the gap between cosmological simulations and small-scale MBH dynamics, we introduce the “Massive Black Hole Assembly in Galaxies Informed by Cosmological Simulations” (MAGICS) suite, which combines the realistic MBH and galaxy population from the state-of-the-art cosmological hydrodynamics simulation ASTRID with the idealized galaxy merger simulations. This suite (also referred to as “resimulations” hereafter) directly recovers the high-redshift galaxy merger properties to include both isolated and multiple galaxy mergers, and uses the “full-physics” hydrodynamical subgrid modeling with star formation and various feedback channels. This is the first work to recover not only the collisionless component (e.g. dark matter and stars) in cosmological MBH merger events but also the full hydrodynamics evolution of the merging system. The importance of the latter in gas-rich environments (typical of high-redshift galaxies) has already been pointed out in e.g. Fiacconi et al. (2013), Tamburello et al. (2017), Liao et al. (2023b).
@@ 2 Idealized Galaxy Simulation with MP-Gadget
Figure 1: Top Row: Evolution of a disk galaxy with the “SH03” model in MP-Gadget after 500 Myrs. 2D projected gas density (left), density-weighted temperature (middle) and 2D projected density of newly-formed stars (right). Middle Row: Evolution of the same disk galaxy run with the ASTRID subgrid model in MP-Gadget after 500 Myrs. Bottom Row: evolution of the same disk galaxy as the middle row run with the ASTRID subgrid model, but with $50\%$ of the gas put into the gas halo component.
@@ 2.1 The subgrid physics model for galaxy formation
The dynamics of the BHs are modeled with a sub-grid dynamical friction model (Tremmel et al., 2015; Chen et al., 2021) to replace the original implementation that directly repositions the BHs to the minimum local potential. This model provides an improved treatment for calculating BH trajectories and velocities. Two BHs merge if their separation is within two times the spatial resolution $2\epsilon_{\rm g,BH}$, once their kinetic energy is dissipated by dynamical friction, and they are gravitationally bound to the local potential. In the ASTRID simulation, we use a separate mass tracer $M_{\rm dyn}$ to reduce the noisy gravitational forces (dynamical heating) acting on the small seed mass black holes during the force calculations of BHs (gravity and dynamical friction). When a new BH is seeded, $M_{\rm dyn}$ is set to $M_{\rm dyn,seed}=10^{7}{h^{-1}M_{\odot}}$, which is about $1.5M_{\rm DM}$. $M_{\rm dyn}$ is kept at its seeding value until $M_{\rm BH}>M_{\rm dyn,seed}$. After that, $M_{\rm dyn}$ grows following the BH mass accretion.
@@ 2.2 Code Validation
Figure 2: Comparison between the gas surface density profile (top panel), gas disk height (middle panel), and star-formation rate (SFR) surface density between the three runs. The lines with light colors show the gas properties in the initial conditions, and the lines with dark colors show the properties after 1 Gyr of evolution. The “SH03” run (purple) and the “Astrid Model” run (green) share the same IC (light blue). The “Astrid Model+Gas Halo” run (orange) has $50\%$ gas in the disk and $50\%$ gas in the halo for the initial condition (light orange).
Figure 3: Evolution of the total star formation rate in the three validation runs. Without wind and AGN feedback, the “SH03” run has the most (up to $\sim 4$ times higher) star formation throughout the simulation. The model with gaseous halo displays more efficient star formation compared to the no-halo run because more central gas is pushed to further distances by wind feedback in the disk-only model.
@@ 2.2.1 Disk Galaxy with Star Formation
We first validate the physical models and implementations in MP-Gadget in the context of an isolated disk galaxy. We set up the initial condition (IC) following the Agora code comparison suite (Kim et al., 2014, 2016), using the MakeDisk IC generator Springel et al. (2005), with parameters matching with Table 3 of Kim et al. (2016), except that we use a disk gas fraction of 0.4 to match our high-redshift application. We model the multiphase interstellar medium following the prescription of Springel & Hernquist (2003) that incorporates gas cooling, star formation, and SN thermal feedback. (this run is denoted as “SH03”). In this first validation step, we did not match the full ASTRID physical models: we did not include the wind feedback from TypeII supernova and the influence of molecular hydrogen on star-formation; we also did not include BHs in this run, so there is no BH accretion and feedback. We run the simulation for $1\,{\rm Gyr}$ and observe a stable disk throughout the simulation.
@@ 2.2.2 Disk Galaxy with ASTRID Models
To match the physical models used in the ASTRID simulation, as a further validation of the resimulation subgrid modeling we include a BH with an initial mass of $4\times 10^{6}\,M_{\odot}$ and turn on all the ASTRID subgrid physics models described in Section 2. Compared with the vanilla SH03 modeling shown in the previous section, adding wind and AGN feedback is expected to remove the dense gas at the galaxy center and regulate the star formation in the disk (e.g. Di Matteo et al., 2005b; Okamoto et al., 2010; Weinberger et al., 2017). We use two different initial conditions for this “Astrid-Model” run: we first keep the same IC as the “SH03” run in the previous section, with only the addition of a central BH. Then, to match more closely with the high-redshift gas environments where only a small fraction of the total gas in the halo is star-forming and in disk structure ($\sim 10-40\%$, see e.g. Table 4 in Appendix A), we use a modified version of MakeDisk to put $50\%$ of the total gas into the gaseous halo component, following the method laid out in Su et al. (2019). We set up the gas halo in thermal equilibrium following a $\beta$ profile with $\beta=0.66$ and $R_{c}/R_{s}=0.5$. We adiabatically relax the IC for $250\,{\rm Myrs}$ until it becomes stable, before turning on other subgrid models.
@@
In Figure 1 we show comparisons of projected gas properties and newly formed stars in the SH03 run, the “Astrid Model” run with all gas in the disk, and the “Astrid Model” run with a gas halo (“Astrid Model+Gas Halo”). With feedback models turned on, we see a significant drop in central star formation (the right column), and the disk gas can cool to much lower temperatures (middle column). We find that the supernova wind feedback is more efficient at removing gas from the galaxy center and thus lowering the central SFR, compared with AGN feedback. Figure 2 shows the comparisons between the gas profiles in the three runs, along with their corresponding ICs (note that the “SH03” run and the “Astrid Model” run share the same IC so it is only shown once by the light blue curve). We see that the disk height and a high star-formation rate are maintained throughout the “SH03” run. The “Astrid Model+Gas Halo” run has a flattened star-formation surface density and gas surface density after a Gyr of evolution.
@@
Finally, Figure 3 shows the total SFR in the isolated galaxies in the three runs during 1 Gyr. The “SH03” run has the most (up to $\sim 4$ times higher) star formation throughout the simulation. The model with a gaseous halo displays more efficient star formation compared to the no-halo run because more central gas is pushed to further distances by wind feedback in the disk-only model. In the “Astrid Model+Gas Halo” run we observe a continuous inflow of cold gas clumps, which sustains the star formation in the disk for a longer time. These clumps can also impact the dynamics of seed MBHs. We will use the “Astrid Model+Gas Halo” modeling in all of the resimulations in this work.
@@ 3 High-Redshift MBH Mergers
Figure 4: The $z=6$ MBH merger population in Astrid. From left to right: the mass distribution of the more massive MBH among the pair (first panel); the distribution of the mass ratio between the two merging MBHs (second panel); the number of MBHs in the central region ($<3{\rm kpc}$ from the galaxy center) of the remnant galaxy (third panel); the seeding redshift of the more massive BH in the pair (fourth panel). The background green histogram shows the entire merger population, and we also show two sub-populations: the blue histogram is the ”isolated” mergers with no third MBH coming into the central region of the host galaxy before $z=6$, and the red histogram shows the ”complex” mergers with multiple MBH in the host galaxy center already at $z=6$. The numbers overlaid on each plot label where the resimulated systems lie within each distribution (only the x-values are meaningful, and the y-values are randomly taken).
@@ 3.1 $z\sim 6$ merger population in ASTRID
In Figure 4, we show the MBH properties of all mergers in this redshift bin, including the primary mass, mass ratio, the total number of MBHs in the remnant galaxy, and the seeding redshift of the primary MBH. We subdivide the merger population into three groups, according to the total number of MBHs in the remnant galaxy at $z=6$ (we take this as a proxy for the complexity of the host environment). The first group is the “isolated mergers” which have no new infalling structures interacting with the merging MBHs at least until $z=6$ (blue population in Figure 4). This group consisted of the majority of high-z mergers (1531/2107), with the merger remnant being the only MBH in its host galaxy. Within these “isolated mergers”, however, a significant fraction will have new infalling MBHs/galaxies soon after the ASTRID merger at $z=6$ (i.e. we found new structures on the outskirts of the remnant halo at $z=6$). Although these new infalls will not impact MBHs that already merged in ASTRID by $z=6$, they may interfere with the binary formation in the high-resolution resimulations. Hence, we treat these systems separately (shown as the orange bar in the third panel of Figure 4) when setting up resimulations.
@@
$M_{\rm DM}$ $M_{\rm gas}$ $M_{\rm star}$ $M_{\rm BH,dyn}$ $\epsilon_{\rm DM}$ $\epsilon_{\rm gas}$ $\epsilon_{\rm star}$ $\epsilon_{\rm BH}$
@@
The second group consists of “multiple MBH” systems (the red population), where at $z=6$, the galaxy remnant has $1-2$ other MBHs besides the merging pair. This means that the orbits of the pair go through more complex interactions already before $z=6$ with other galaxies/MBHs. From Figure 4, we see that about $25\%$ (542 out of 2107) of $z=6$ MBH pairs reside in these multiple galaxy interaction environments. This highlights the importance of considering multiple MBH interactions when modeling the dynamics of MBHs even in such high-redshift mergers. Finally, we leave out a group of the most complex (with $>3$ MBHs in the remnant galaxy) merger systems for this work. This group makes up a very small fraction of the total merger population (34 out of 2107). In future works, it is still worthwhile to study these systems, as they often reside in the high-density peaks of the Universe and may trace the merger events during the formation of the first quasars.
@@
The right two panels of Figure 5 show the power-law index of the density profiles and the densities measured at the ASTRID resolution limit (similar to the method used in Chen et al., 2022). The gravitational potential of high-redshift galaxies is dominated by the dark matter halo above kpc scales. In most cases, the dark matter density exceeds the stellar density by a factor of $\sim 10$. However, these galaxies are gas-rich and mergers can also trigger a phase of rapid star formation. Therefore, as we will also show later, the stellar densities are subjected to growth by a factor of $\sim 10$ over the timescale of a few hundred Myrs, and can dominate over dark matter on sub-kpc scales.
@@ 4.1 Initial Conditions
Figure 7 shows the comparisons between the ASTRID merging system and the resimulation initial conditions generated as described above. In the left and middle columns, we show visualizations of the host galaxies and gas environments, with matched color scales between Astrid and the resimulation (brightness corresponds to the density, and gas is color-coded by temperature with blue corresponding to cold gas). The resimulation IC resembles the morphologies, stellar densities, and gas temperature of the original system, but shows significant improvement in mass and spatial resolution. The improvement in resolution can be seen more clearly from the left column, where we show the density profiles of the original halos and the resimulation halos. The detailed properties of all components in each recovered galaxy are shown in Appendix A.

### v3

@@ Abstract
Keywords: Massive Black Holes, Gravitational Waves, Galaxies, Hydrodynamical Simulations, High-Redshift.
@@ 1 Introduction
Currently, our most accurate understanding of the MBH binary population and its relation with galaxy evolution comes from cosmological simulations (see e.g. Amaro-Seoane et al., 2023, for an overview). Cosmological simulations self-consistently follow the coevolution of MBHs with host galaxies, and contain rich information about the environments where MBH interactions and mergers take place (e.g. Salcido et al., 2016; Kelley et al., 2017; Tremmel et al., 2017; Volonteri et al., 2020; Katz et al., 2020; Chen et al., 2022). These environments include a wide range of scenarios from isolated dwarf galaxies to infalling satellites of a massive central galaxy in a cluster. However, realistic and large-volume modeling comes at the cost of limited resolutions, and at best they can follow MBH dynamics to $\sim{\rm kpc}$ scales. These simulations are also subject to the simplistic subgrid-seeding mechanism that only considers the higher end of the MBH seed mass (out of these, Dubois et al., 2014; Tremmel et al., 2015, probes relatively low MBH seeds at high-redshifts). To compensate for the resolution limit in large-volume simulations, cosmological “zoom-in" simulations are also applied to study the sinking behavior of specific merging systems (e.g. Pfister et al., 2019; Bortolas et al., 2020), but they are also subjected to high-computational cost and low flexibilities in configuration and subgrid models. Only a few galaxies and the MBHs can be studied at a time.
@@
To truly bridge the gap between cosmological simulations and small-scale MBH dynamics, we introduce the “Massive Black Hole Assembly in Galaxies Informed by Cosmological Simulations" (MAGICS) suite, which combines the realistic MBH and galaxy population from the state-of-the-art cosmological hydrodynamics simulation ASTRID with the idealized galaxy merger simulations. This suite (also referred to as “resimulations" hereafter) directly recovers the high-redshift galaxy merger properties to include both isolated and multiple galaxy mergers, and uses the “full-physics" hydrodynamical subgrid modeling with star formation and various feedback channels. This is the first work to recover not only the collisionless component (e.g. dark matter and stars) in cosmological MBH merger events but also the full hydrodynamics evolution of the merging system. The importance of the latter in gas-rich environments (typical of high-redshift galaxies) has already been pointed out in e.g. Fiacconi et al. (2013), Tamburello et al. (2017), Liao et al. (2023b).
@@ 2 Idealized Galaxy Simulation with MP-Gadget
Figure 1: Top Row: Evolution of a disk galaxy with the “SH03" model in MP-Gadget after 500 Myrs. 2D projected gas density (left), density-weighted temperature (middle) and 2D projected density of newly-formed stars (right). Middle Row: Evolution of the same disk galaxy run with the ASTRID subgrid model in MP-Gadget after 500 Myrs. Bottom Row: evolution of the same disk galaxy as the middle row run with the ASTRID subgrid model, but with $50\%$ of the gas put into the gas halo component.
@@ 2.1 The subgrid physics model for galaxy formation
The dynamics of the BHs are modeled with a sub-grid dynamical friction model (Tremmel et al., 2015; Chen et al., 2021) in both ASTRID and the resimulations. This model provides an improved treatment for calculating BH trajectories and velocities. Two BHs merge if their separation is within two times the spatial resolution $2\epsilon_{\rm g,BH}$ (this is $\sim 500\,{\rm pc}$ in ASTRID at $z=6$, and $20\,{\rm pc}$ in the resimulations), once their kinetic energy is dissipated by dynamical friction, and they are gravitationally bound to the local potential. We note that since we numerically merge the MBHs at $20\,{\rm pc}$ in the resimulations, our modeling ends before the MBH pairs become a bound binary. Therefore we do not attempt to model binary formation and binary hardening process (except in Section 5.5). In the ASTRID simulation, we use a separate mass tracer $M_{\rm dyn}$ to reduce the noisy gravitational forces (dynamical heating) acting on the small seed mass black holes during the force calculations of BHs (gravity and dynamical friction). When a new BH is seeded, $M_{\rm dyn}$ is set to $M_{\rm dyn,seed}=10^{7}{h^{-1}M_{\odot}}$, which is about $1.5M_{\rm DM}$. $M_{\rm dyn}$ is kept at its seeding value until $M_{\rm BH}>M_{\rm dyn,seed}$. After that, $M_{\rm dyn}$ grows following the BH mass accretion.
@@ 2.2 Code Validation
Figure 2: Comparison between the gas surface density profile (top panel), gas disk height (middle panel), and star-formation rate (SFR) surface density between the three runs. The lines with light colors show the gas properties in the initial conditions, and the lines with dark colors show the properties after 1 Gyr of evolution. The “SH03" run (purple) and the “Astrid Model" run (green) share the same IC (light blue). The “Astrid Model+Gas Halo" run (orange) has $50\%$ gas in the disk and $50\%$ gas in the halo for the initial condition (light orange).
Figure 3: Evolution of the total star formation rate in the three validation runs. Without wind and AGN feedback, the “SH03" run has the most (up to $\sim 4$ times higher) star formation throughout the simulation.
@@ 2.2.1 Disk Galaxy with Star Formation
We first validate the physical models and implementations in MP-Gadget in the context of an isolated disk galaxy. We set up the initial condition (IC) following the Agora code comparison suite (Kim et al., 2014, 2016), using the MakeDisk IC generator Springel et al. (2005), with parameters matching with Table 3 of Kim et al. (2016), except that we use a disk gas fraction of 0.4 to match our high-redshift application. We model the multiphase interstellar medium following the prescription of Springel & Hernquist (2003) that incorporates gas cooling, star formation, and SN thermal feedback. (this run is denoted as “SH03"). In this first validation step, we did not match the full ASTRID physical models: we did not include the wind feedback from TypeII supernova and the influence of molecular hydrogen on star-formation; we also did not include BHs in this run, so there is no BH accretion and feedback. We run the simulation for $1\,{\rm Gyr}$ and observe a stable disk throughout the simulation.
@@ 2.2.2 Disk Galaxy with ASTRID Models
To match the physical models used in the ASTRID simulation, as a further validation of the resimulation subgrid modeling we include a BH with an initial mass of $4\times 10^{6}\,M_{\odot}$ and turn on all the ASTRID subgrid physics models described in Section 2. Compared with the vanilla SH03 modeling shown in the previous section, adding wind and AGN feedback is expected to remove the dense gas at the galaxy center and regulate the star formation in the disk (e.g. Di Matteo et al., 2005b; Okamoto et al., 2010; Weinberger et al., 2017). We use two different initial conditions for this “Astrid-Model" run: we first keep the same IC as the “SH03" run in the previous section, with only the addition of a central BH. Then, to match more closely with the high-redshift gas environments where only a small fraction of the total gas in the halo is star-forming and in disk structure ($\sim 10-40\%$, see e.g. Table 4 in Appendix A), we use a modified version of MakeDisk to put $50\%$ of the total gas into the gaseous halo component, following the method laid out in Su et al. (2019). We set up the gas halo in thermal equilibrium following a $\beta$ profile with $\beta=0.66$ and $R_{c}/R_{s}=0.5$. We adiabatically relax the IC for $250\,{\rm Myrs}$ until it becomes stable, before turning on other subgrid models.
@@
In Figure 1 we show comparisons of projected gas properties and newly formed stars in the SH03 run, the “Astrid Model" run with all gas in the disk, and the “Astrid Model" run with a gas halo (“Astrid Model+Gas Halo"). With feedback models turned on, we see a significant drop in central star formation (the right column), and the disk gas can cool to much lower temperatures (middle column). We find that the supernova wind feedback is more efficient at removing gas from the galaxy center and thus lowering the central SFR, compared with AGN feedback. Figure 2 shows the comparisons between the gas profiles in the three runs, along with their corresponding ICs (note that the “SH03" run and the “Astrid Model" run share the same IC so it is only shown once by the light blue curve). We see that the disk height and a high star-formation rate are maintained throughout the “SH03" run. The “Astrid Model+Gas Halo" run has a flattened star-formation surface density and gas surface density after a Gyr of evolution.
@@
Finally, Figure 3 shows the total SFR in the isolated galaxies in the three runs during 1 Gyr. The “SH03" run has the most (up to $\sim 4$ times higher) star formation throughout the simulation. The model with a gaseous halo displays more efficient star formation compared to the no-halo run because more central gas is pushed to further distances by wind feedback in the disk-only model. In the “Astrid Model+Gas Halo" run we observe a continuous inflow of cold gas clumps, which sustains the star formation in the disk for a longer time. These clumps can also impact the dynamics of seed MBHs. We will use the “Astrid Model+Gas Halo" modeling in all of the resimulations in this work.
@@ 3 High-Redshift MBH Mergers
Figure 4: The $z=6$ MBH merger population in Astrid. From left to right: the mass distribution of the more massive MBH among the pair (first panel); the distribution of the mass ratio between the two merging MBHs (second panel); the number of MBHs in the central region ($<3{\rm kpc}$ from the galaxy center) of the remnant galaxy (third panel); the seeding redshift of the more massive BH in the pair (fourth panel). The background green histogram shows the entire merger population, and we also show two sub-populations: the blue histogram is the "isolated" mergers with no third MBH coming into the central region of the host galaxy before $z=6$, and the red histogram shows the "complex" mergers with multiple MBH in the host galaxy center already at $z=6$. The numbers overlaid on each plot label where the resimulated systems lie within each distribution (only the x-values are meaningful, and the y-values are randomly taken).
@@ 3.1 $z\sim 6$ merger population in ASTRID
In Figure 4, we show the MBH properties of all mergers in this redshift bin, including the primary mass, mass ratio, the total number of MBHs in the remnant galaxy, and the seeding redshift of the primary MBH. We subdivide the merger population into three groups, according to the total number of MBHs in the remnant galaxy at $z=6$ (we take this as a proxy for the complexity of the host environment). The first group is the “isolated mergers" which have no new infalling structures interacting with the merging MBHs at least until $z=6$ (blue population in Figure 4). This group consisted of the majority of high-z mergers (1531/2107), with the merger remnant being the only MBH in its host galaxy. Within these “isolated mergers", however, a significant fraction will have new infalling MBHs/galaxies soon after the ASTRID merger at $z=6$ (i.e. we found new structures on the outskirts of the remnant halo at $z=6$). Although these new infalls will not impact MBHs that already merged in ASTRID by $z=6$, they may interfere with the binary formation in the high-resolution resimulations. Hence, we treat these systems separately (shown as the orange bar in the third panel of Figure 4) when setting up resimulations.
@@
[…]
@@
The second group consists of “multiple MBH" systems (the red population), where at $z=6$, the galaxy remnant has $1-2$ other MBHs besides the merging pair. This means that the orbits of the pair go through more complex interactions already before $z=6$ with other galaxies/MBHs. From Figure 4, we see that about $25\%$ (542 out of 2107) of $z=6$ MBH pairs reside in these multiple galaxy interaction environments. This highlights the importance of considering multiple MBH interactions when modeling the dynamics of MBHs even in such high-redshift mergers. Finally, we leave out a group of the most complex (with $>3$ MBHs in the remnant galaxy) merger systems for this work. This group makes up a very small fraction of the total merger population (34 out of 2107). In future works, it is still worthwhile to study these systems, as they often reside in the high-density peaks of the Universe and may trace the merger events during the formation of the first quasars.
@@
The right two panels of Figure 5 show the power-law index of the density profiles and the densities measured at the ASTRID resolution limit (similar to the method used in Chen et al., 2022). Specifically, we assume that below a scale $r_{\rm ext}$ close to the resolution limit $2.8\times\epsilon_{g}=0.85\,{\rm kpc}$, the stellar density profile follows a single power-law $\rho\propto r^{-\gamma}$. To measure the value of $\gamma$, we take the measured density from 10 bins just above $r_{\rm ext}$, and fit it to the power-law profile. The gravitational potential of high-redshift galaxies is dominated by the dark matter halo above kpc scales. In most cases, the dark matter density exceeds the stellar density by a factor of $\sim 10$. However, these galaxies are gas-rich and mergers can also trigger a phase of rapid star formation. Therefore, as we will also show later, the stellar densities are subjected to growth by a factor of $\sim 10$ over the timescale of a few hundred Myrs, and can dominate over dark matter on sub-kpc scales.
@@ 4.1 Initial Conditions
Figure 7 shows the comparisons between the ASTRID merging system and the resimulation initial conditions generated as described above. In the left and middle columns, we show visualizations of the host galaxies and gas environments, with matched color scales between Astrid and the resimulation (brightness corresponds to the density, and gas is color-coded by temperature with blue corresponding to cold gas). The resimulation IC resembles the morphologies, stellar densities, and gas temperature of the original system, but shows significant improvement in mass and spatial resolution. The improvement in resolution can be seen more clearly from the left column, where we show the density profiles of the original halos and the resimulation halos (measured after the adiabatic relaxation, just before we start the resimulations). The detailed properties of all components in each recovered galaxy are shown in Appendix A.